An automobile radiator contains 16L of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should be drained and replaced with pure antifreeze so that the mixture will be 50% antifreeze?
Answer by stanbon(75887)

An automobile radiator contains 16L of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should be drained and replaced with pure antifreeze so that the mixture will be 50% antifreeze?
-----
Equation:
alcohol - alcohol + alcohol = alcohol
----
0.30*16 - 0.30x + 1.00x = 0.50*16
-------------------
30*16 - 30x + 100x = 50*16
70x = 20*16
x = (2/7)16 = 4.57 L (amt. to be removed and replaced)
=====================
Cheers,
Stan H.